OpenStudy (shelby1290):

WILL GIVE MEDAL TO BEST ANSWER!! Factor and show your work a)8x^2+10x+3 b) 6m^2-3m-3 c)2a^2-11a+12

Directrix (directrix):

I'll do the first one and save the other two. 8x^2+10x+3 is of the form ax^2 + bx + c a = 8 b = 10 c = 3 To get started, look for two numbers that multiply to a*c and add to b. Multiply to 8*3 = 24 and add to 10. 6 and 4 multiply to 24 and add to 10. "Bust" the b 8x^2+10x+3 8 x^2 + 6x + 4x + 3 = 2x* ( 4x + 3) + 1*(4x + 3) = (4x + 3) * ( 2x + 1)

OpenStudy (shelby1290):

@Directrix i'm not sure if i did part C correctly and i don't know where i might have messed up

OpenStudy (shelby1290):

2a^2-11a+12 2 X 12=24 8 X 3=24 8+3=11 =2a^2+8a+3a+12 =(2a^2+8a)(3a+12) =2a(a+4)+3(a+4) =(a+4)(2a+3)

Directrix (directrix):

c) 2a^2-11a+12 ---------------- I think there is a sign error but am not sure. Let me look.

OpenStudy (shelby1290):

Ohhh i think i figured it out... It's supposed to be -8 X -3 =24 and (-8)+(-3)=11

Directrix (directrix):

2a^2-11a+12 2 X 12=24 The two numbers have to multiply to 24 and add to -11 They would be -8 and -3.

OpenStudy (shelby1290):

okay well for part B, i can't figure out anything else that equals to -18 and adds up to -3 6m^2-3m-3 6 X -3 = -18

Directrix (directrix):

It's supposed to be -8 X -3 =24 and (-8)+(-3)= - 11 Correct if you may that negative 11

Directrix (directrix):

Before going to that problem, this current problem where =(a+4)(2a+3) appears is the answer needs correcting.

OpenStudy (shelby1290):

unless the answer is just -6 and 3

OpenStudy (shelby1290):

yes, i've already fixed the question on my paper

Directrix (directrix):

2a^2-8a-3a+12 2a^2 - 8a -3a + 12 2a(a -4) -3 ( a - 4) (2a - 3) ( a - 4) -->

OpenStudy (shelby1290):

part c) =2a^2-8a-3a+12 =(2a-8a)(-3a+12) =(2a(a-4)+3(a-4) =(a-4)(2a-3)

Directrix (directrix):

Okay.

OpenStudy (shelby1290):

now how about for part b? are the numbers -6 and 3?

Directrix (directrix):

b) 6m^2-3m-3 Multiply to -18 and add to -3

OpenStudy (shelby1290):

-6 x3=-18 -6+3=-3

Directrix (directrix):

Oh, first divide through by 3 because that is a common factor. 6m^2-3m-3 = 3*( 2m^2 -m -1) Focus on this: ( 2m^2 -m -1) What are two numbers that multiply to -2 and add to -1 ?

OpenStudy (shelby1290):

-2 X 1=-2 -2 + 1= -1

Directrix (directrix):

We will put back the initial 3 after factoring this: ( 2m^2 -m -1) 2 m^2 -2m + 1m - 1 = 2m(m -1) + 1(m - 1) = (m - 1) (2m +1) The final answer: 6m^2-3m-3 = 3*(m - 1) (2m +1)

Directrix (directrix):

Is that what you got?

OpenStudy (shelby1290):

yes, i did! thank you

Directrix (directrix):

You are welcome.